Find the absolute value of -2|7-3y|-6=-14
Found 2 solutions by ptth176, Alan3354:
Answer by ptth176(50)

-2|7-3y|-6=-14
=>|7-3y|=4
=>7-3y=4 or 7-3y=-4
=>y=1 or y=11/3

Answer by Alan3354(69443)

Find the absolute value of -2|7-3y|-6=-14
-----------------
That makes no sense.
Maybe you want to solve for y ???
-------------
-2|7-3y|-6=-14
Add 6
-2|7-3y| = -8
Divide by -2
|7-3y| = 4
-------------
Case 1:
7-3y = 4
-3y = -3
y = 1
-----------------
Case 2:
-(7-3y) = 4
7+3y = -4
3y = -11
y = -11/3